TufStuff, Inc., sells a wide range of drums, bins, boxes, and other containers that are used in the chemical industry. One of the company’s products is a heavy-duty corrosion-resistant metal drum, called the WVD drum, used to store toxic wastes. Production is constrained by the capacity of an automated welding machine that is used to make precision welds. A total of 2,100 hours of welding time is available annually on the machine. Because each drum requires 0.4 hours of welding machine time, annual production is limited to 5,125 drums. At present, the welding machine is used exclusively to make the WVD drums. The accounting department has provided the following financial data concerning the WVD drums:

WVD Drums
Selling price per drum $ 169.00
Cost per drum:
Direct materials $52.10
Direct labor ($23 per hour) 4.60
Manufacturing overhead 7.50
Selling and administrative expense 30.80 95.00
Margin per drum $ 74.00

Management believes 6,125 WVD drums could be sold each year if the company had sufficient manufacturing capacity. As an alternative to adding another welding machine, management has considered buying additional drums from an outside supplier. Harcor Industries, Inc., a supplier of quality products, would be able to provide up to 4,125 WVD-type drums per year at a price of $153 per drum, which TufStuff would resell to its customers at its normal selling price after appropriate relabeling.

Megan Flores, TufStuff’s production manager, has suggested the company could make better use of the welding machine by manufacturing bike frames, which would require only 0.5 hours of welding machine time per frame and yet sell for far more than the drums. Megan believes that TufStuff could sell up to 1,700 bike frames per year to bike manufacturers at a price of $264 each. The accounting department has provided the following data concerning the proposed new product:

Bike Frames
Selling price per frame $ 264.00
Cost per frame:
Direct materials $102.40
Direct labor ($18 per hour) 36.80
Manufacturing overhead 41.00
Selling and administrative expense 51.80 232.00
Margin per frame $ 32.00

The bike frames could be produced with existing equipment and personnel. Manufacturing overhead is allocated to products on the basis of direct labor-hours. Most of the manufacturing overhead consists of fixed common costs such as rent on the factory building, but some of it is variable. The variable manufacturing overhead has been estimated at $1.35 per WVD drum and $1.90 per bike frame. The variable manufacturing overhead cost would not be incurred on drums acquired from the outside supplier.

Selling and administrative expenses are allocated to products on the basis of revenues. Almost all of the selling and administrative expenses are fixed common costs, but it has been estimated that variable selling and administrative expenses amount to $0.75 per WVD drum whether made or purchased and would be $1.80 per bike frame.

All of the company’s employees—direct and indirect—are paid for full 40.00-hour work weeks and the company has a policy of laying off workers only in major recessions.

As soon as your analysis was shown to the top management team at TufStuff, several managers got into an argument concerning how direct labor costs should be treated when making this decision. One manager argued that direct labor is always treated as a variable cost in textbooks and in practice and has always been considered a variable cost at TufStuff. After all, “direct” means you can directly trace the cost to products. “If direct labor is not a variable cost, what is?” Another manager argued just as strenuously that direct labor should be considered a fixed cost at TufStuff. No one had been laid off in over a decade, and for all practical purposes, everyone at the plant is on a monthly salary. Everyone classified as direct labor works a regular 40.00-hour workweek and overtime has not been necessary since the company adopted Lean Production techniques. Whether the welding machine is used to make drums or frames, the total payroll would be exactly the same. There is enough slack, in the form of idle time, to accommodate any increase in total direct labor time that the bike frames would require.

Required:

1. Would you be comfortable relying on the financial data provided by the accounting department for making decisions related to the WVD drums and bike frames?

2. Compute the contribution margin per unit for [assume direct labor is a fixed cost]

3. Compute the contribution margin per welding hour for [assume direct labor is a fixed cost]

4. Assuming direct labor is a fixed cost:

a. Determine the number of WVD drums (if any) that should be purchased and the number of WVD drums and/or bike frames (if any) that should be manufactured.

b. What is the increase (decrease) in net operating income that would result from this plan over current operations?

5. Compute the contribution margin per unit for [assume direct labor is a variable cost]

6. Compute the contribution margin per welding hour for [assume direct labor is a variable cost]

7. Assuming direct labor is a variable cost:

a. Determine the number of WVD drums (if any) that should be purchased and the number of WVD drums and/or bike frames (if any) that should be manufactured. [Assume direct labor is a variable cost]

b. What is the increase (decrease) in net operating income that would result from this plan over current operations?
Answer #1
1 No
Margin computed by the accounting department is not useful for decision making since the cost includes allicated fixed common costs that are irrelavant to decision making
2 Assuming direct labor is a fixed cost,
Contribution margin per unit:
Purchased
WVD drums
Manufactured
WVD drums
Manufactured
Bike
frames
Selling price a 169 169 264
Less; Variable costs
Materials 153 52.1 102.4
Direct labor 0 0 0
Variable manufacturing overhead 0 1.35 1.9
Variable selling and administrative 0.75 0.75 1.8
Total variable cost b 153.75 54.2 106.1
Contribution margin per unit (a-b) 15.25 114.8 157.9
3 Contribution margin per welding hour:
Manufactured
WVD drums
Manufactured
Bike
frames
Contribution margin per unit a 114.8 157.9
Welding hours per unit b 0.4 0.5
Contribution margin per welding hour (a/b) 287 315.8
4
a. Ranking of alternatives based on contribution margin per hour:
Manufactured
WVD drums
Manufactured
Bike
frames
Contribution margin per welding hour (a/b) 287 315.8
Rank 2 1
Preference should be given to manufacturing bike frames
Total hours available=2100 hours
Demand for Bike frames=1700 frames
Welding hours per unit=0.50
Welding hours used for bike frames=1700*0.50=850 hours
Balance welding hours available=2100-850=1250 hours
Welding hours per unit for WVD drums=0.40
Drums can be produced=1250/0.40=3125 Drums
Demand for WDV drums=6125 drums
Rest of the drums can be purchased from outside
Drums to be purchased=6125-3125=3000 drums
Purchased Manufactured
WVD drums 3000 3125
Bike frames 0 1700
b. Increase in net operating income:
Units Contribution
margin per
unit
Total
contribution
Manufactured
WVD drums
3125 114.8 358750
Purchased
WVD drums
3000 15.25 45750
Bike frames produced 1700 157.9 268430
Total 672930
Less: Current contribution margin 5125 114.8 588350
Increase in net operating income 84580
5 Assuming direct labor is a fixed cost,
Contribution margin per unit:
Purchased
WVD drums
Manufactured
WVD drums
Manufactured
Bike
frames
Selling price a 169 169 264
Less; Variable costs
Materials 153 52.1 102.4
Direct labor 0 4.6 36.8
Variable manufacturing overhead 0 1.35 1.9
Variable selling and administrative 0.75 0.75 1.8
Total variable cost b 153.75 58.8 142.9
Contribution margin per unit (a-b) 15.25 110.2 121.1
6 Contribution margin per welding hour:
Manufactured
WVD drums
Manufactured
Bike
frames
Contribution margin per unit a 110.2 121.1
Welding hours per unit b 0.4 0.5
Contribution margin per welding hour (a/b) 275.5 242.2
7
a. Ranking of alternatives based on contribution margin per hour:
Manufactured
WVD drums
Manufactured
Bike
frames
Contribution margin per welding hour (a/b) 275.5 242.2
Rank 1 2
Preference should be given to manufacturing WVD drums
Total hours available=2100 hours
Welding hours per unit for WVD drums=0.40
Drums can be produced=2100/0.40=5250 Drums
Demand for WDV drums=6125 drums
Rest of the drums can be purchased from outside
Drums to be purchased=6125-5250=875 drums
Purchased Manufactured
WVD drums 875 5250
Bike frames 0 0
b. Increase in net operating income:
Units Contribution
margin per
unit
Total
contribution
Manufactured
WVD drums
5250 110.2 578550
Purchased
WVD drums
875 15.25 13344
Bike frames produced 0 121.1 0
Total 591894
Less: Current contribution margin 5125 110.2 564775
Increase in net operating income 27119